OpenStudy (kuvyojhmoob):

A rectangular plot of ground having dimensions 29 ft by 30 ft is surrounded by a walk of uniform width. If the area of the walk is 536 ft2, what is its width?

OpenStudy (campbell_st):

let the uniform width be x  so the area of the whole thing plot and path is \[(2x + 30)(2x + 29) = 29 \times 30 + 536\] you have a quadratic equation that will allow you to solve for x \[4x^2 + 118x + 870 = 1406\] hope it helps